# Prompt-Based Bias Calibration for Better Zero/Few-Shot Learning of Language Models

### v1

@@ 5 Experiments
We conduct extensive experiments on 8 English datasets, including sentiment analysis and topic classification. They consist of 5 sentence-level datasets potentially impacted by common token bias: AGNews Zhang et al. (2015), DBPedia Lehmann et al. (2015), TREC Voorhees and Tice (2000), Subj Pang and Lee (2004), SST-5 Socher et al. (2013) and 3 aspect-level sentiment analysis datasets likely subject to association bias: Restaurant and Laptop reviews from SemEval 2014 Task Pontiki et al. (2014), Twitter Dong et al. (2014). For these aspect-level datasets, the task is to predict sentiments associated with the marked aspects in each sentence. More details are in Appendix A and Table 7.
@@ 5.1 Evaluation Protocol
We evaluate the effectiveness of our intrinsic-bias calibration method on enhancing Masked LMs zero/few-shot learning performance with 4 prompt learning methods: in-context learning and prompt-based fine-tuning, both with and without demonstration. We follow the prompt-based fine-tuning and demonstration method of Gao et al. (2021).
We conduct calibration with 5 different seeds, and for the few-shot setting, we randomly sample 5 different groups of training and validation sets ( samples per class). We report the mean and standard deviation of LM performance. For the 5 sentence-level classification tasks, we use accuracy as the metric. For the 3 aspect-level classification tasks, because of the imbalance in test set, we use weighted F for a balanced evaluation. Details of calibration and prompt learning are in Appendix A.
We present our main results using RoBERTa-large, and for few-shot setting. Results of using RoBERTa-base, , and different prompt templates are in Appendix B.
@@ 5.2 Main Results
The output representations of <mask> token for label word predictions are visualized by t-SNE van der Maaten and Hinton (2008) in Figure 4. On the left, samples from the two categories are almost mixed together, indicating that the original LM tends to bias toward one class prediction. In contrast, the right visualization demonstrates improved separability after One-batch Calibration(§ 3.3), which explains the significant performance enhancement achieved by our intrinsic-bias calibration method.
In-context lrn no demo | Prompt FT no demo | |||
UnSel. | Sel. | UnSel. | Sel. | |
AGNews | ||||
DBPedia | ||||
TREC | ||||
Subj | ||||
@@
In-context lrn no demo | Prompt FT no demo | |||
WLM + BLM | BLM | WLM + BLM | BLM | |
AGNews | ||||
DBPedia | ||||
TREC | ||||
Subj | ||||
SST-5 | ||||
Laptop | ||||
Restaurant | ||||
@@
5.3 Update Entire LM vs. Only Bias Parameters in Calibration
In Table 4, we evaluate the impact of updating entire LM (WLM + BLM) during calibration on downstream task performance, as compared to only updating bias parameters (BLM) in calibration. The optimal learning rate for updating entire LM in calibration is smaller as shown in Appendix A Table 6. Results of In-context lrn/Prompt FT with demo are in Appendix B (Table 11). For in-context learning, the LM with only BLM updates in calibration achieves better overall performance compared to the LM with entire parameter updates, most likely attributed to better preserved language modeling abilities. For prompt-based fine-tuning, two differently calibrated LMs demonstrate comparable performance, as the impact of entire-parameter calibration on the modeling ability is mitigated through task-specific fine-tuning. Considering the significant saving in memory and computation, we recommend only updating BLM in calibration.
@@ 5.4 Analysis
Evaluate language modeling abilities after calibration. Our method calibrates the intrinsic bias associated with a set of task-specific label words. In this section, we explore the impact of updating LM for task-specific bias calibration on general language modeling abilities. Specifically, we take the LM calibrated for one task and evaluate its performance on the other tasks. The results are shown in Figure 5. In general, intrinsic bias calibration for one task has a minimal adverse effect on other tasks’ modeling (no more than 2% degradation) due to the light model updates, while remarkably enhancing LM performance on that specific task. Notably, there is consistent performance increase at bottom right, as these tasks are all sentiment classification sharing or including same label words.444For aspect-level datasets, better improvement is on the diagonals (task-specific calibration), indicating our method mitigates the impact of association bias (Appendix A).
Prompt FT no demo | Prompt FT with demo | |||
NoisyTune | IntrCal | NoisyTune | IntrCal | |
AGNews | ||||
DBPedia | ||||
TREC | ||||
Subj | ||||
SST-5 | ||||
Laptop | ||||
Restaurant | ||||
Average | 79.3 | 80.0 |
Is the improvement in prompt-based fine-tuning simply attributed to NoisyTune? Wu et al. (2022) demonstrates that adding noise to pre-trained LMs benefits conventional fine-tuning on downstream tasks. To validate that the gains in prompt-based fine-tuning with our method are not solely a result of perturbing LM parameters, we conduct comparison experiments by adding noise (with the most suitable intensity suggested in Wu et al., 2022) to the pre-trained LM before initiating prompt-based fine-tuning. Table 5 illustrates that, while NoisyTune proves effective in better fine-tuning pre-trained LMs on downstream tasks (as compared to NoCal in Table 2), our method consistently surpasses NoisyTune, confirming the efficacy of intrinsic bias calibration in enhancing LM performance.
@@ 7 Limitations
Second, we utilize RoBERTa (encoder) models for classification tasks, as they have demonstrated competitive performance Gao et al. (2021); Logan IV et al. (2022). However, the relatively small size of those Masked LMs (355M parameters for RoBERTa-large) could be the ultimate limitation to their capabilities. Given the proliferation of large-scale decoder LMs and their accomplishments in tackling more challenging tasks Thoppilan et al. (2022); Chowdhery et al. (2023); Touvron et al. (2023), we anticipate extending our method to large decoder models and validating the applicability of our findings. Furthermore, we expect to expand the scope of tasks to include regression problems (e.g., sentiment score prediction) leveraging KL divergence to measure disparities in continuous probability distributions, aiming to address bias-related challenges across diverse scenarios.
8 Ethics Statement and Broader Impact
@@ Appendix A Experimental Details
Hyper-parameters. In calibration stage, we shuffle the null-input prompts and conduct gradient descent on BLM (or WLM + BLM as comparative experiment) with 5 different seeds to account for calibration variance. There are two main hyper-parameters for calibration: (1) batch size ; (2) calibration learning rate . We conduct grid search on and , and obtain the best settings: and as shown in Table 6.
Calibrated LMs are applied in downstream tasks with prompt-learning methods. We use the same hyper-parameters as Gao et al. (2021) for prompt learning. We evaluate on each task’s original test set, except for AGNews and DBPedia, where we randomly sample 2000 test examples.
We use PyTorch Paszke et al. (2019) and public HuggingFace Transformers library Wolf et al. (2020), and conduct all the experiments with one NVIDIA V100 GPU in Google Colab.
Calibration | Prompt FT (downstream) | ||
@@ Appendix B Additional Results
In-context lrn no demo | In-context lrn with demo | Prompt FT no demo | Prompt FT with demo | |||||||||
NoCal | OutCal | IntrCal | NoCal | OutCal | IntrCal | NoCal | OutCal | IntrCal | NoCal | OutCal | IntrCal | |
AGNews | ||||||||||||
DBPedia | ||||||||||||
TREC | ||||||||||||
Subj | ||||||||||||
SST-5 | ||||||||||||
Laptop | ||||||||||||
Restaurant | ||||||||||||
@@
Task | Prompt Templates |
In-context lrn with demo | Prompt FT with demo | |||
WLM + BLM | BLM | WLM + BLM | BLM | |
AGNews | ||||
DBPedia | ||||
TREC | ||||
Subj | ||||
SST-5 | ||||
Laptop | ||||
Restaurant | ||||

### v2

@@ 5 Experiments
We conduct extensive experiments on 8 English datasets, including sentiment analysis and topic classification.333 We mainly focus on single-sentence tasks, which aligns with the use of single-sentence null inputs for calibration. The alignment may enhance calibration effectiveness. We also experiment on sentence-pair tasks in Appendix B.3 Table 18 and demonstrate better performance after calibration. They consist of 5 sentence-level datasets potentially impacted by common token bias: AGNews Zhang et al. (2015), DBPedia Lehmann et al. (2015), TREC Voorhees and Tice (2000), Subj Pang and Lee (2004), SST-5 Socher et al. (2013) and 3 aspect-level sentiment analysis datasets likely subject to association bias: Restaurant and Laptop reviews from SemEval 2014 Task Pontiki et al. (2014), Twitter Dong et al. (2014). For aspect-level datasets, the task is to predict sentiments associated with the marked aspects in each sentence. More details are in Appendix A Table 7.
@@ 5.1 Evaluation Protocol
We evaluate the effectiveness of our intrinsic-bias calibration method on enhancing Masked LMs zero/few-shot learning performance with 4 prompt-based learning methods: in-context learning and prompt-based fine-tuning, both with and without demonstration. We follow the prompt-based fine-tuning and demonstration method of Gao et al. (2021). Besides Masked LMs, we also validate the effectiveness of our method on two decoder LMs: GPT-2 XL (1.5B) Radford et al. (2019) and Llama-2 (7B) Touvron et al. (2023) in Appendix B.2.
We conduct calibration with 5 different seeds, and for the few-shot setting, we randomly sample 5 different groups of training and validation sets ( samples per class). We report the mean and standard deviation of LM performance. For the 5 sentence-level classification tasks, we use accuracy as the metric. For the 3 aspect-level classification tasks, because of the imbalance in test set, we use weighted F1 for a balanced evaluation. Details of calibration and prompt-based learning are in Appendix A.
@@ 5.2 Main Results
In addition to NoCal and OutCal, we compare our results with those reproduced from NoisyTune Wu et al. (2022), NSP-BERT Sun et al. (2022) and Perplection Lu et al. (2023), as detailed in Appendix B.1 (Table 8, 9). The superior performance further validates the effectiveness of our method.
@@
In-context lrn no demo | Prompt FT no demo | ||||
UnSel. | Sel. | UnSel. | Sel. | ||
AGNews | |||||
DBPedia | |||||
TREC | |||||
Subj | |||||
@@
The output representations of <mask> token for label word predictions are visualized by t-SNE in Figure 3. On the left, samples from the two categories are almost mixed together, indicating that the original LM tends to bias toward one class prediction. In contrast, the right visualization demonstrates improved separability after One-batch Calibration(§ 3.3), which explains the significant performance enhancement achieved by our intrinsic-bias calibration method.
5.3 Update Entire LM vs. Only Bias Parameters in Calibration
In Table 4, we evaluate the impact of updating entire LM (WLM + BLM) during calibration on downstream task performance, as compared to only updating bias parameters (BLM). The optimal learning rate for updating entire LM is smaller (Appendix A Table 6). For in-context learning, the LM with only BLM updates in calibration achieves better overall performance compared to the LM with entire parameter updates, most likely attributed to better preserved language modeling abilities (Appendix B.3 Table 16). For prompt-based fine-tuning, two differently calibrated LMs demonstrate comparable performance, as the impact of entire-parameter calibration on the modeling ability is mitigated through task-specific fine-tuning. Considering the significant saving in memory and computation, we recommend only updating BLM in calibration.
In-context lrn no demo | Prompt FT no demo | ||||
WLM + BLM | BLM | WLM + BLM | BLM | ||
AGNews | |||||
DBPedia | |||||
TREC | |||||
Subj | |||||
SST-5 | |||||
Laptop | |||||
Restaurant | |||||
@@ 5.4 Analysis
How does intrinsic bias calibration impact downstream tasks? Our method calibrates the intrinsic bias associated with a set of task-specific label words. In this section, we explore the impact of updating LM for task-specific bias calibration on other downstream task performance. Specifically, we take the LM calibrated for one task and evaluate its performance on the other tasks as shown in Figure 4. In general, intrinsic bias calibration for one task has a minimal adverse effect on other tasks’ performance (no more than 2% degradation) because of the light model updates, while remarkably enhancing LM performance on that specific task. Notably, there is consistent performance increase at bottom right, as these tasks are all sentiment classification sharing or including same label words.444For aspect-level datasets, better improvement is on the diagonals (task-specific calibration), indicating our method mitigates the impact of association bias (Appendix A).
How does intrinsic bias calibration impact language modeling abilities? We employ pseudo-perplexity Salazar et al. (2020) to evaluate language modeling for Masked LM. Following each task-specific intrinsic bias calibration, we measure pseudo-perplexity and compare the results with original RoBERTa on WikiText-2, WikiText-103 Merity et al. (2017), and LAMBADA dataset Paperno et al. (2016). As shown in Table 5, language modeling abilities are largely preserved after calibration due to the minimal updates to the model.
WT-2 | WT-103 | LAMBADA | |
Original RoBERTa | |||
+ calibration | |||
for_AGNews | |||
for_DBPedia | |||
for_TREC | |||
for_Subj | |||
for_SST-5 | |||
for_Laptop | |||
for_Restaurant | |||
for_Twitter |
@@ 6 Conclusion
Second, we utilize RoBERTa (encoder) models for classification tasks, as encoder models may more effectively encode task-specific patterns for discriminative tasks compared to some generative LMs Gao et al. (2021); Li et al. (2023b), as shown in Table 12. However, the relatively small size of those Masked LMs (355M parameters for RoBERTa-large) could be the ultimate limitation to their capabilities. Given the proliferation of large-scale generative (decoder) LMs and their accomplishments in tackling more challenging tasks Thoppilan et al. (2022); Chowdhery et al. (2023); Touvron et al. (2023), we anticipate extending our method to large decoder models and validating the applicability of our findings. Furthermore, we expect to expand the scope of tasks to include regression problems (e.g., sentiment score prediction) leveraging KL divergence to measure disparities in continuous probability distributions, aiming to address bias-related challenges across diverse scenarios.
Ethics Statement and Broader Impact
@@
Acknowledgments
This work was supported in part by the Center for Co-Design of Cognitive Systems (CoCoSys), a Semiconductor Research Corporation (SRC) and DARPA-sponsored JUMP 2.0 center.
@@ Appendix A Experimental Details
Null-meaning inputs generation with GPT-4. The version of GPT-4 used in our experiment is gpt-4-0613. We observe that GPT-4 could generate repetitive null-meaning inputs. To avoid overrepresentation of certain null inputs which might impact the diversity and introduce bias to the null-input set, we adopt an iterative approach. In each iteration, GPT-4 generates 500 null-meaning inputs, and duplicates are removed. This process continues until we obtain 1000 distinct null-meaning inputs, which takes 3 iterations in our experiment.
Null-meaning inputs for One-batch Calibration. For zero-shot downstream tasks, since only one batch of null-meaning inputs is required for calibration in our early-stopping criterion (§ 3.3), we select the from , where is batch size. We prioritize these samples as our observations show that null-meaning inputs with higher exhibit higher attention scores between the null input and <mask>, as demonstrated in Figure 5. This indicates more effective conveyance of the "null" information to the placeholder <mask>, which could facilitate LM deciphering the "null" patterns of the prompts and benefit calibration.
Hyper-parameters. In calibration stage, we shuffle the null-input prompts and conduct gradient descent on BLM (or WLM + BLM as comparative experiment) with 5 different seeds to account for calibration variance. There are two main hyper-parameters for calibration: (1) batch size ; (2) calibration learning rate . We conduct grid search on and from to , and obtain the best settings: and as shown in Table 6.
Calibrated LMs are applied in downstream tasks with prompt-based learning methods. We use the same hyper-parameters as Gao et al. (2021) for prompt-based learning. We evaluate on each task’s original test set, except for AGNews and DBPedia, where we randomly sample 800 test examples.
We use PyTorch Paszke et al. (2019) and public HuggingFace Transformers library Wolf et al. (2020). RoBERTa related experiments are conducted on a single NVIDIA V100 GPU, while GPT-2 and Llama-2 experiments are conducted on one A100 GPU in Google Colab.
Calibration () | Prompt FT (downstream) | ||
@@ Appendix B Additional Results
B.1 Performance Comparison with NSP-BERT, Perplection and NoisyTune
We additionally choose NSP-BERT Sun et al. (2022) and Perplection Lu et al. (2023) as in-context learning comparison baselines and NoisyTune Wu et al. (2022) as prompt-base fine-tuning comparison baseline. NSP-BERT constructs potential answers using each label word and predict Next Sentence Prediction (NSP) probability between the input and each answer. Perplection proposes perplexity-based selection method for prompt-based zero-shot learning. NoisyTune demonstrates that adding noise to pre-trained LMs benefits fine-tuning on downstream tasks. We re-implement their methods with the same settings as ours for fair comparisons. As shown in Table 8 and Table 9, our method achieves superior results in almost all datasets.
Furthermore, our method consistently outperforms NoisyTune, demonstrating that the gains in prompt-based fine-tuning with our method are not solely a result of perturbing LM parameters. This confirms the efficacy of intrinsic bias calibration in enhancing LM performance.
Zero-shot in-context learning | |||
NSP-BERT | Perplection | IntrCal | |
AGNews | 54.5 | ||
DBPedia | 61.8 | ||
TREC | 32.4 | ||
Subj | 62.7 | ||
SST-5 | 37.5 | ||
Laptop | 59.6 | ||
Restaurant | 72.8 | ||
51.7 | |||
Average | 54.0 |
Prompt FT no demo | Prompt FT with demo | ||||
NoisyTune | IntrCal | NoisyTune | IntrCal | ||
AGNews | |||||
DBPedia | |||||
TREC | |||||
Subj | |||||
SST-5 | |||||
Laptop | |||||
Restaurant | |||||
Average | 79.3 | 80.0 |
B.2 Effectiveness on Decoder LMs
We validate the effectiveness of intrinsic bias calibration in enhancing prompt-based learning performance on GPT-2 XL (1.5B) and Llama-2 (7B). The same hyper-parameters from Appendix A and prompt templates from Table 7 are used for bias calibration. For GPT-2, we only update the bias parameters during calibration, whereas for Llama-2, we update the entire model since it does not have bias parameters. We conduct zero-shot and two-shot in-context learning experiments across the eight classification datasets, comparing original (Orig.) LM and calibrated (Calib.) LM. The performance comparisons are shown in Table 10 (GPT-2) and Table 11 (Llama-2). Calibrated LMs demonstrate significant performance improvement compared to original pre-trained LMs.
Zero-shot | Two-shot | ||||
Orig. LM | Calib. LM | Orig. LM | Calib. LM | ||
AGNews | |||||
DBPedia | |||||
TREC | |||||
Subj | |||||
SST-5 | |||||
Laptop | |||||
Restaurant | |||||
Average | 45.7 | 58.6 |
Zero-shot | Two-shot | ||||
Orig. LM | Calib. LM | Orig. LM | Calib. LM | ||
AGNews | |||||
DBPedia | |||||
TREC | |||||
Subj | |||||
SST-5 | |||||
Laptop | |||||
Restaurant | |||||
Average | 50.6 | 62.0 |
In Table 12, we compare the performance of RoBERTa-large (355M) with GPT-2 XL (1.5B) and Llama-2 (7B) in zero-shot learning on classification tasks, using their original pre-trained models. RoBERTa outperforms the other models on more datasets, and achieves better computing efficiency due to its smaller model size. Encoder LMs could be more effective and efficient for classification tasks for several reasons: (i) The bidirectional architecture of encoder LMs enables them to capture task-specific patterns more effectively by attending to both left and right context, compared to the unidirectional nature of decoder LMs. (ii) Classification tasks prioritize accurate label prediction over the generation of diverse and human-like text. Besides, the label spaces in classification are significantly more constrained than the whole vocabulary used in generative applications, which may restrict the effectiveness of decoder LMs Li et al. (2023b). (iii) The relative small size of encoder models facilitates efficiently combining prompting with label-supervised fine-tuning for classification tasks Liu et al. (2023), which further enhances performance, as demonstrated in Table 2.
B.3 Other Experiments
We briefly summarize the contents of each table and figure below that presents other additional results.
Table 14 contains results for performance using RoBERTa-base model.
Table 15 contains results for performance of few-shot learning.
Table 16 contains results for pseudo-perplexity comparisons between updating entire LM and only updating bias parameters in calibration.
Table 17 contains results for performance comparisons between updating entire LM and only updating bias parameters in calibration.
Table 18 contains results for performance of sentence-pair datasets.
Table 19 contains results for variance of probability distribution across labels before and after calibration.
RoBERTa-large | GPT-2 XL | Llama-2 | |
(355M) | (1.5B) | (7B) | |
AGNews | 47.0 | ||
DBPedia | 58.2 | ||
TREC | 42.0 | ||
Subj | 50.8 | ||
SST-5 | 33.2 | ||
Laptop | 54.6 | ||
Restaurant | 68.6 | ||
25.5 | |||
Average | 44.3 |
Task | Prompt Templates |
In-context lrn no demo | In-context lrn with demo | Prompt FT no demo | Prompt FT with demo | ||||||||||||
NoCal | OutCal | IntrCal | NoCal | OutCal | IntrCal | NoCal | OutCal | IntrCal | NoCal | OutCal | IntrCal | ||||
AGNews | |||||||||||||||
DBPedia | |||||||||||||||
TREC | |||||||||||||||
Subj | |||||||||||||||
SST-5 | |||||||||||||||
Laptop | |||||||||||||||
Restaurant | |||||||||||||||
@@ B.3 Other Experiments
Model | Datasets | |||||||
WikiText-2 | WikiText-103 | LAMBADA | ||||||
Original RoBERTa | ||||||||
[…]
+ calibration | WLM + BLM | BLM | WLM + BLM | BLM | WLM + BLM | BLM | ||
[…]
for_AGNews | ||||||||
for_DBPedia | ||||||||
for_TREC | ||||||||
for_Subj | ||||||||
for_SST-5 | ||||||||
for_Laptop | ||||||||
for_Restaurant | ||||||||
for_Twitter |
ICL with demo | Prompt FT with demo | ||||
WLM + BLM | BLM | WLM + BLM | BLM | ||
AGNews | |||||
DBPedia | |||||
TREC | |||||
Subj | |||||
SST-5 | |||||
Laptop | |||||
Restaurant | |||||
@@
In-context lrn no demo | Prompt FT no demo | ||||
NoCal | IntrCal | NoCal | IntrCal | ||
MNLI | |||||
SNLI | |||||
MRPC | |||||
QQP |